# The Impact of Copyrighted Material on Large Language Models: A Norwegian Perspective

### v1

@@ Abstract
The use of copyrighted materials in training generative language models raises critical legal and ethical questions. This paper presents a framework for and the results of empirically assessing the impact of copyrighted materials on the performance of large language models (LLMs) for Norwegian. We found that both books and newspapers contribute positively when the models are evaluated on a diverse set of Norwegian benchmarks, while fiction works possibly lead to decreased performance. Our experiments could inform the creation of a compensation scheme for authors whose works contribute to AI development.
@@ 1 Introduction
Generative language models have fundamentally reshaped the landscape of natural language processing (NLP), enabling the development of systems that can generate and interact with human language at an unprecedented level. This includes Norwegian, for which several large language models (LLMs) have been trained and published in the recent years using different architectures and licensing choices Kummervold et al. (2021); Kutuzov et al. (2021); Samuel et al. (2023); Liu et al. (2024).
@@
However, the vast quantities of data required for training these models often include copyrighted materials, presenting novel challenges related to intellectual property rights and compensation. These challenges have led to numerous lawsuits across jurisdictions, fundamentally questioning the legitimacy of training models on copyrighted data without explicit permissions from content creators (Panettieri, 2024; Madigan, 2024; Weisenberger et al., 2024).
@@
The first wave of lawsuits emerged shortly after the public release of advanced generative AI models. Content creators, including authors, visual artists, and musicians, began to express concerns about the unauthorized use of their work in training datasets. Multiple class-action lawsuits were filed in the United States, accusing prominent AI companies such as OpenAI and Meta Platforms of infringing on copyright laws by using copyrighted materials without obtaining explicit permissions. The authors argued that the unauthorized use of their works without any form of compensation or recognition undermines their intellectual property rights and jeopardizes their ability to earn a living from their creative endeavors. In Europe, a coalition of news publishers has taken legal action against Google and Meta Platforms, arguing that the use of journalistic content in training models without fair remuneration constitutes a breach of copyright and undermines the sustainability of high-quality journalism. Likewise, Norwegian rightsholder organizations representing publishing houses across the country, contacted the government in late 2023 expressing their concerns over the use of their material in training generative language models and demanding some sort of compensation were their contents to be used in the training of generative language models.
@@
In this context, this paper describes a first attempt at empirically evaluating the impact of copyrighted content in the training of LLMs for Norwegian. We introduce a set of carefully curated datasets that are later used in the training of foundational, domain-tuned, and instruction-tuned models. We establish the proper training conditions to be able to compare models trained on the different datasets. A newly created benchmarking suite is used to evaluate the performance of each individual model and make the comparison meaningful. As a collaborative effort among several institutions, we hope the results of our investigations set the basis to guide policymaking and proper compensation schemes for authors and rightsholders in Norway.
@@ 3 Data Collection
At the core of our research sits the need for guidelines and well-designed compensation mechanisms in Norway. Hence, we had to abide by Norwegian law (Åndsverkloven) for the definition of which works are protected under copyright in the country and what we are permitted to do with said works (which might differ from other jurisdictions). In line with provisions that allow research on language technology and data mining, and with the consent of the Norwegian rightsholder organizations, this study primarily relied on material legally deposited at the National Library of Norway. Materials not deposited were therefore omitted. As the institution has digitized almost all books and around 85% of the newspapers ever published in the country Nasjonalbiblioteket (2024), we believe this set of works is representative enough to draw valid conclusions from its study.
We collaborated with the National Library of Norway and the rightsholder organizations to gain access to protected materials for the building of our datasets. Others sources were also added via specific agreements with third party organizations such as the Norwegian Broadcasting Corporation (NRK), the TV channel TV2, and the newspaper conglomerates Amedia and Schibsted. Publicly available text collections like Wikipedia, datasets from the HPLT de Gibert et al. (2024) and CulturaX Nguyen et al. (2024) projects, and code in different programming languages from Lozhkov et al. (2024) were also added with the objective of setting up a realistic training scenario where using Internet crawled sources is commonplace (see Figure 1). This mixture of data allowed us to evaluate the impact of high-quality copyrighted material versus non-copyrighted alternatives. The models trained on the copyrighted materials will not be made publicly available for further use and only serve the purpose of this study.
@@ 3.1 Core Datasets
Figure 1: Treemap with the number of words contributed by each source in the extended dataset.
We followed the recipe of the Norwegian Colossal Corpus (NCC) by Kummervold et al. (2022), adapting and updating it with new up-to-date contents, re-OCRing some materials, enriching their metadata, and ensuring a uniform format and functionality across datasets. The preparation involved cleaning, deduplication, metadata tagging, and language balancing to maintain consistent representation of Norwegian, preventing other languages from overshadowing it. The corpus was divided into two main datasets: a base dataset excluding copyrighted materials [^1] Except for newspapers that fall under the Language Technology Use (Språkteknologiformål), as they were already included in other datasets such as NCC., and an extended dataset that included all collected texts, thus including all of base (see Table 1).
@@
We decided to include texts from other Scandinavian (Swedish, Danish, and Icelandic) and English sources to boost the performance of the resulting language models. To ensure that languages other than Norwegian, and primarily coming via Internet crawling, were balanced, we adapted the perplexity-based sampling strategy from De la Rosa et al. (2022) to maintain a high quality in the selected data. Instead of sampling a fixed number of documents, parameters for a Gaussian curve were calculated from 500,000-1M random documents per source, utilizing Wikipedia-based Kneser-Ney language models from Wenzek et al. (2019) and Conneau et al. (2020) [^2] Built with KenLM Heafield (2011).. We also modified the perplexity calculation to account for normalized text. These parameters then guided dataset sub-sampling to target ratios per language, reducing foreign language content while maintaining quality.
Moreover, we also built three custom perplexity models for specific Norwegian domains that proved too divergent from Wikipedia: books, newspapers, and government documents. These perplexity models were used to score each document in the datasets. Based on their perplexity scores, the documents were further divided into three segments corresponding to their quartile distribution. Documents with scores below the first quartile were classified as “good”, those between $Q_{1}$ and the third quartile as “medium”, and those above $Q_{3}$ were considered “bad”. To ensure robust training, the documents in each segment were randomized. It is also important to notice than in order to maintain the language distributions for foreign languages with respect to the amount of Norwegian texts, the total number of documents in foreign languages in the extended dataset is consequently higher and slightly different (due to the sampling strategies) than that of base; we keep the same ratios.
@@ 3.2 Domain Specific Subsets
Table 2: Number of documents and words in each copyrighted subset of the ‘extended’ dataset.
The copyrighted materials present in the extended dataset were further divided into groups. These subsets were carefully designed to test the effect of specific features of the copyrighted content. As illustrated in Table 2, we created subsets for books vs newspapers, fiction vs factual content, and original Norwegian vs translations.
@@ 3.3 Instruction-tuning Datasets
To align the model more closely with human objectives and assess whether instruction tuning with limited high-quality data can enhance the performance of our pre-trained models across various tasks, we built upon prior work and collected nearly 5,000 instructions annotated by research assistants formatted as (instruction, input, output) triplets. In this format, instruction refers to the directive provided by humans for the model, input is an optional field containing task-related information, and output denotes the desired response that follows the given instruction.
@@
[…]
@@
[…]
@@ 4 Model Training
    To measure the overall impact of copyrighted material and its effect in realistic scenarios, we conducted pre-training on the base and extended datasets, both from scratch and ‘warm-started’: using the existing weights of the pre-trained model Mistral 7B v0.1333https://huggingface.co/mistralai/Mistral-7B-v0.1. These four core models were trained on the same amount of data (64,000 steps of 4 million sub-word tokens each) using identical sets of hyperparameters.
@@
    To further isolate the effect of different ablations of the copyrighted materials, we continuously fine tuned the model trained on base from scratch for an extra 10,000 steps on each of the 9 domain specific subsets.
@@
Overall, we trained 17 models (7 billion parameters each) using a total of 270,000 GPU-hours. Model training specifications are shown in Table 3.3. The infrastructure for training included the LUMI supercomputer, Idunn cluster, and Google TPUs through the Tensor Research Cloud program. [^4] To assess the deviation introduced by differences in training infrastructures and platforms across the participating institutions, each team trained a control model with 1.5B parameters based on the Llama 2 architecture. The training setups were identical, utilizing the base dataset. After comparing the validation loss curves from each team, we found that the curves were almost the same, with a deviation of less than $0.05$ in terms of the final convergence validation loss. Besides, we trained two tokenizers with the base and extended datasets separately, both with the same vocabulary size of $32,768$. After an initial test of the fertility [^5] Fertility expresses the fragmentation rate of a tokenizer and is #$\#tokens/\#words$ in one corpus. of the tokenizers, we found the difference between them was only $0.0013$. Therefore, we decided to use the same tokenizer trained with the base dataset for all the models.
@@ 5 Evaluation Framework
Evaluating generative LLMs is far from a solved problem, in particular for Norwegian, where there were few existing resources at the outset of our work. We compiled a set of 28 common NLP tasks, encompassing a range of different metrics to assess the performance of each of the models. These tasks can be grouped into 9 higher level skills:
@@
    Sentiment Analysis, here defined as binary polarity classification on both the sentence- and document-level, based on the existing NoReC datasets of professional reviews Velldal et al. (2018); Øvrelid et al. (2020).
@@
    Fairness & Truthfulness. Fairness refers to the absence of bias in the predictions and outputs of a model. Evaluating fairness ensures that the model does not favor or discriminate against particular groups based on attributes like race, gender, or ethnicity. This skill was evaluated using a newly-created dataset, [^6] https://huggingface.co/datasets/mimir-project/mimir-bias which covers a wide range of bias types, including race, religion, gender, geography, occupation, age etc. Truthfulness involves the accuracy and reliability of the information produced by the model, ensuring it generates factual and verifiable content. This skill was evaluated using the newly-created NorTruthfulQA dataset which adapts the English TruthfulQA dataset Lin et al. (2021) to Norwegian and assesses whether a model is truthful in selecting and generating answers to questions that involve common human misconceptions. [^7] https://huggingface.co/datasets/ltg/nortruthfulqa_mc and https://huggingface.co/datasets/ltg/nortruthfulqa_gen
@@
    Reading Comprehension, which measures the ability of a model to understand and interpret text. It involves answering questions about a given passage, summarizing content, or explaining the meaning of specific phrases or sentences. This skill estimates how well the model grasps the context and details in the text. It was evaluated using the existing extractive question-answering NorQUAD dataset Ivanova et al. (2023).
@@
    World Knowledge, which assesses the extent of factual information a language model has about the world. This includes historical events, geographical data, scientific facts, cultural knowledge, and more. The model should correctly answer questions or provide information based on real-world knowledge. This skill was evaluated using the newly-created NorOpenBookQA [^8] https://huggingface.co/datasets/ltg/noropenbookqa dataset, which adapts the English OpenBookQA dataset Mihaylov et al. (2018) to Norwegian and the NRK-Quiz-QA [^9] https://huggingface.co/datasets/ltg/nrk dataset, consisting of a collection of online quizzes on a number of different topics.
@@
    Commonsense Reasoning, which involves the ability of a model to make logical inferences based on everyday knowledge and understanding of the world. The model should reason about situations that require practical, everyday knowledge that people take for granted. This skill was evaluated using the newly-created NorCommonsenseQA dataset [^10] https://huggingface.co/datasets/ltg/norcommonsenseqa of multiple-choice commonsense question answer-pairs which adapts the corresponding English CommonsenseQA dataset Talmor et al. (2019) to Norwegian.
@@
    Summarization, which measures the ability of a model to condense longer pieces of text into shorter, coherent summaries that capture the main points. This skill is crucial for applications where users need a quick understanding of large volumes of information, such as news articles or research papers. It was evaluated using the MediaFutures datase. [^12] https://huggingface.co/datasets/SamiaT/NorSumm
@@
    Translation, which assesses how accurately a language model can convert text from one language to another while preserving the meaning, tone, and context. It was evaluated using the existing Tatoeba dataset Tiedemann (2020).
@@
[A complete description of the evaluation framework is available in a forthcoming publication which is currently under review.] All the newly created datasets will be made publicly available upon paper acceptance. We follow the standard in-context learning evaluation design for pretrained decoder-only language models (e.g., Brown et al., 2020; Touvron et al., 2023), which includes zero-shot and few-shot evaluation.
In this paper, for the sake of simplicity, we selected the most common metrics per task and aggregated scores using a simple cumulative sum per higher-level skill. In order to aggregate results into an overall score, with the caveats of aggregating metrics of different nature, scores were extracted for the best available {0, 1, 4, 16}-shot configuration for each task and the best score for each of the prompts (we formulated five different prompts for each task, but the evaluation is based on the best performing prompt for each of them). Metrics were normalized to exhibit the same higher-is-better behavior in a range of 0 to 100.
base (warm)extended (warm)base + translated booksbase + original books + newspapersbase + original booksbase + nonfiction books + newspapersbase + nonfiction booksbase + fiction booksbase + books + newspapersbase + newspapersbase + booksbaseextended441.83413.98427.30440.35435.64408.20427.40440.97427.87436.81409.54482.98480.28Sentiment AnalysisFairness & TruthfulnessReading ComprehensionWorld KnowledgeCommonsense ReasoningNorwegian LanguageSummarizationTranslationVariation & Readability

### v2

@@ Abstract
The use of copyrighted materials in training language models raises critical legal and ethical questions. This paper presents a framework for and the results of empirically assessing the impact of publisher-controlled copyrighted corpora on the performance of generative large language models (LLMs) for Norwegian. When evaluated on a diverse set of tasks, we found that adding both books and newspapers to the data mixture of LLMs tend to improve their performance, while the addition of fiction works seems to be detrimental. Our experiments could inform the creation of a compensation scheme for authors whose works contribute to AI development.
@@ 1 Introduction
Generative language models have radically reshaped the landscape of natural language processing (NLP), enabling the development of systems that can generate and interact with human language at an unprecedented level. This includes Norwegian, for which several large language models (LLMs) have been trained and published in the recent years using different architectures and licensing choices Kummervold et al. (2021); Kutuzov et al. (2021); Samuel et al. (2023, 2025); Liu et al. (2024).
@@
However, the vast quantities of data required for training these models often include copyrighted materials, presenting novel challenges related to intellectual property rights and compensation. Additionally, prior research has highlighted significant concerns about dataset composition and quality in large-scale web-crawled datasets, emphasizing the need for more responsible data curation practices (Kreutzer et al., 2022; Artetxe et al., 2022; Penedo et al., 2024). Together, these challenges have led to numerous lawsuits across jurisdictions, fundamentally questioning the legitimacy of training models on copyrighted data without explicit permissions from content creators (Panettieri, 2024; Madigan, 2024; Weisenberger et al., 2024) [^1] See Gervais et al. (2024) for an in-depth introduction on how LLMs are being interpreted in the legal domain..
@@
The first wave of lawsuits emerged shortly after the public release of advanced generative AI models (see Appendix A). Content creators, including authors, visual artists, and musicians, began to express concerns about the unauthorized use of their work in training datasets. Multiple class-action lawsuits were filed in the United States, accusing prominent AI companies such as OpenAI and Meta Platforms of infringing on copyright laws by using copyrighted materials without obtaining explicit permissions. The authors argued that the unauthorized use of their works without any form of compensation or recognition undermines their intellectual property rights and jeopardizes their ability to earn a living from their creative endeavors. In Europe, a coalition of news publishers has taken legal action against Google and Meta Platforms, arguing that the use of journalistic content in training models without fair remuneration constitutes a breach of copyright and undermines the sustainability of high-quality journalism. Likewise, Norwegian rights-holder organizations representing publishing houses across the country, contacted the government in late 2023 expressing their concerns over the use of their material in training generative language models and demanding some sort of compensation were their contents to be used in the training of generative language models. As a result, the Ministry of Culture and Equality (Kultur- og likestillingsdepartemente) instructed the National Library to create a data-driven report they could use in order to make informed decisions in the elaboration of a compensation scheme for the authors. Led by the National Library of Norway, a consortium was formed together with the University of Oslo and the Norwegian University of Science and Technology under the umbrella of the so-called Mímir Project [^2] A name chosen after a figure in Norse mythology renowned for his knowledge and wisdom..
@@
In this context, and under the umbrella of Mímir, this paper describes a first attempt at empirically evaluating the impact of copyrighted content in the training of LLMs for Norwegian. We introduce a set of carefully curated datasets that are later used in the training of foundational, domain-tuned, and instruction-tuned models. We establish the proper training conditions to be able to compare models trained on the different datasets. A newly created benchmarking suite is used to evaluate the performance of each individual model and make the comparison meaningful. As a collaborative effort among several institutions, the results of our investigations set the basis to guide policymaking and proper compensation schemes for authors and right-holders in Norway (Nasjonalbiblioteket, 2024).
@@ 3 Data Collection
With the objective of setting up a realistic training scenario where using Internet crawled sources is commonplace, we gathered publicly available text collections like Wikipedia, datasets from the HPLT de Gibert et al. (2024) and CulturaX Nguyen et al. (2024) projects, code in different programming languages from Lozhkov et al. (2024), governmental reports and publications published under open licenses, and books and newspapers articles in the public domain.
We then collaborated with the National Library of Norway and the rights-holder organizations to gain access to protected materials. Through the legal deposit act, the National Library of Norway has digitized almost all books in Norwegian and around 85% of the newspapers ever published in the country Nasjonalbiblioteket (2024). Where the quality of the digitized material was not enough (e.g., due to OCR processing), or was not been legally deposited (e.g., paywalled content), specific agreements were put in place to obtain the material from third party organizations such as the Norwegian Broadcasting Corporation (NRK), the TV channel TV2, and the newspaper conglomerates Amedia and Schibsted. In line with provisions that allow research on language technology and data mining (Åndsverkloven), and with the consent of the Norwegian right-holders, this study primarily relied on material legally deposited at, or under agreement with, the National Library of Norway. Specifically, we focus our study on the collection of publisher-controlled books and newspapers articles.
@@ 3.1 Core Datasets
Figure 1: Treemap with the final number of words (comma separated) contributed by each source after cleaning and deduplication.
This mixture of data (see Figure 1 and Appendix C) allowed us to evaluate the impact of high-quality publisher-controlled copyright-protected corpora versus other sources commonly available on the Internet. The models trained on the copyrighted materials will not be made publicly available for further use and only serve the purpose of this study.
We followed the recipe of the Norwegian Colossal Corpus (NCC) by Kummervold et al. (2022), adapting and updating it with new up-to-date contents, re-OCRing some materials, enriching their metadata, and ensuring uniform format and functionality across datasets. The preparation involved cleaning, deduplication, metadata tagging, and language balancing to maintain consistent representation of Norwegian, preventing other languages from overshadowing it. The corpus was divided into two main datasets: a base dataset excluding publisher-controlled copyright-protected books and newspapers [^3] Except for newspapers that fall under the Language Technology Use (Språkteknologiformål), as they were already included in other datasets such as NCC., and an extended dataset that included all collected texts, thus including all of base (see Table 1).
@@
We decided to include texts from other Scandinavian (Swedish, Danish, and Icelandic) and English sources to boost the performance of the resulting language models via cross-lingual transfer (Conneau et al., 2020b; Xue et al., 2021). To ensure that languages other than Norwegian, and primarily coming via Internet crawling, were balanced, we adapted the perplexity-based sampling strategy from De la Rosa et al. (2022) to maintain a high quality in the selected data. Instead of sampling a fixed number of documents, parameters for a Gaussian curve were calculated from 500,000-1M random documents per source, utilizing Wikipedia-based Kneser-Ney language models from Wenzek et al. (2019) and Conneau et al. (2020a) [^4] Built with KenLM Heafield (2011).. We also modified the perplexity calculation to account for normalized text. These parameters then guided dataset sub-sampling to target ratios per language, reducing foreign language content while maintaining quality (see Appendix B).
It is also important to notice than in order to maintain the language distributions for foreign languages with respect to the amount of Norwegian texts, the total number of documents in foreign languages in the extended dataset is consequently higher and slightly different (due to the sampling strategies) than that of base; we keep the same ratios (see Appendix C).
@@ 3.2 Domain Specific Subsets
Table 2: Number of documents and words (comma separated) in each subset of the publisher-controlled corpora.
The publisher-controlled copyright-protected materials present in the extended dataset were further divided into groups attending to different criteria. These subsets were carefully designed to test the effect of adding them to the training sets for LLMs. We split the books into fiction vs nonfiction, and original works in Norwegian vs translations. While most books in the collection had metadata information regarding the original language in which a work was written in, genre labels were more scarce. To overcome this limitation, we built a Doc2Vec model [^5] https://huggingface.co/mimir-project/literary-form-classifier Le and Mikolov (2014) that classified fiction vs nonfiction with 98% accuracy and used it to annotate books for which this information was missing. As shown in Table 2, we then built domain specific subsets to investigate 1) the effect of books vs newspapers vs books + newspapers, 2) the effect of factuality by adding only fiction words, only nonfiction works, and nonfiction works + newspapers, and 3) the effect of adding content written originally in Norwegian, such as original books or original books + newspapers, vs translated books.
@@ 3.3 Instruction-tuning Datasets
To align the models more closely with human objectives and assess whether instruction tuning with limited high-quality data can enhance the performance of our pre-trained models across various tasks, we built upon prior work and collected nearly 5,000 instructions annotated by research assistants [^6] https://huggingface.co/datasets/mimir-project/mimir-instruction. The instructions were formatted as (instruction, input, output) triplets, where instruction refers to the directive provided by humans for the model, input is an optional field containing task-related information, and output denotes the desired response that follows the given instruction.
The instruction tuning dataset combines three key categories –Reading Comprehension, Norwegian Culture, and Words and Expressions– with diverse domains to enhance model performance. The domains include Literature, Commonsense, Geography, Language, History, Sports, Entertainment, Food, Politics, Science, Art, Music, and Culture. The variety of the instructions seeks to improve the model’s ability to understand complex texts, provide culturally relevant responses, and handle language nuances, resulting in more versatile, knowledgeable, and context-aware LLMs.
@@
Domain Tuned Models
base + fiction books base 7.5K AMD MI250X
@@
Instruction Fine Tuned Models
base instruct base 14.2 NVIDIA H100
@@ 4 Model Training
    To measure the overall impact of publisher-controlled copyrighted corpora and its effect in realistic scenarios, we conducted pre-training on the base and extended datasets, both from scratch and using the existing weights (warm) of the pre-trained model Mistral 7B v0.1777https://huggingface.co/mistralai/Mistral-7B-v0.1. These four core models were trained on the same amount of data, 64,000 steps of 4 million sub-word tokens each. using identical sets of hyperparameters (see Table 7 in Appendix D). This roughly translates to 3 epochs for the base dataset and 2 for the extended dataset, which according to Muennighoff et al. (2023) is still far from saturating the available data.
@@
    To further isolate the effect of different ablations of the publisher-controlled copyright-protected corpora, we continuously fine tuned the model trained on base from scratch for an extra 10,000 steps on each of the 9 domain specific subsets.
@@
Overall, we trained 17 models (7 billion parameters each) using a total of 270,000 GPU-hours. Model training specifications are shown in Table 3. The infrastructure for training included the LUMI supercomputer, Idunn cluster, and Google TPUs through the Tensor Research Cloud program. [^8] To assess the deviation introduced by differences in training infrastructures and platforms across the participating institutions, each team trained a control model with 1.5B parameters based on the Llama 2 architecture. The training setups were identical, utilizing the base dataset. After comparing the validation loss curves from each team, we found that the curves were almost the same, with a deviation of less than $0.05$ in terms of the final convergence validation loss. Besides, we trained two tokenizers with the base and extended datasets separately, both with the same vocabulary size of $32,768$. After an initial test of the fertility [^9] Fertility expresses the fragmentation rate of a tokenizer and is #$\#tokens/\#words$ in one corpus. of the tokenizers, we found the difference between them was only $0.0013$. Therefore, we decided to use the same tokenizer trained with the base dataset for all the models.
@@ 5 Evaluation Framework
In our empirical evaluation experiments, we utilize NorEval, [^10] https://github.com/ltgoslo/noreval a publicly available framework for evaluating Norwegian generative LLMs built on lm-evaluation-harness Gao et al. (2024). We consider 28 tasks, which test model’s various Norwegian language understanding and generation abilities. NorEval covers both Norwegian language varieties (Bokmål and Nynorsk) and provides a set of 4–6 prompts for each downstream task. The tasks can be grouped into nine higher level skills:
@@
    Sentiment Analysis, here defined as binary polarity classification on both the sentence- and document-level based on the existing NoReC datasets of professional reviews Velldal et al. (2018); Øvrelid et al. (2020).
@@
    Fairness & Truthfulness. Fairness refers to the absence of bias in the predictions and outputs of a model. Evaluating fairness ensures that the model does not favor or discriminate against particular groups based on attributes like race, gender, or ethnicity. This skill was evaluated using a newly-created dataset, [^11] https://huggingface.co/datasets/mimir-project/mimir-bias which covers a wide range of bias types, including race, religion, gender, geography, occupation, age etc. Truthfulness involves the accuracy and reliability of the information produced by the model, ensuring it generates factual and verifiable content. This skill was evaluated using NorTruthfulQA Mikhailov et al. (2025), which assesses whether a model is truthful in selecting and generating answers to questions that involve common human misconceptions. [^12] https://huggingface.co/datasets/ltg/nortruthfulqa_mc and https://huggingface.co/datasets/ltg/nortruthfulqa_gen
@@
    Reading Comprehension, which measures the ability of a model to understand and interpret text. It involves answering questions about a given passage, summarizing content, or explaining the meaning of specific phrases or sentences. This skill estimates how well the model grasps the context and details in the text. It was evaluated using the existing extractive question-answering NorQUAD dataset Ivanova et al. (2023) and multiple-choice question-answering Belebele dataset Bandarkar et al. (2024).
@@
    World Knowledge, which assesses the extent of factual information a language model has about the world. This includes historical events, geographical data, scientific facts, cultural knowledge, and more. The model should correctly answer questions or provide information based on real-world knowledge. This skill was evaluated using the NorOpenBookQA [^13] https://huggingface.co/datasets/ltg/noropenbookqa and NRK-Quiz-QA [^14] https://huggingface.co/datasets/ltg/nrk_quiz_qa by Mikhailov et al. (2025).
@@
    Commonsense Reasoning, which involves the ability of a model to make logical inferences based on everyday knowledge and understanding of the world. The model should reason about situations that require practical, everyday knowledge that people take for granted. This skill was evaluated using NorCommonsenseQA [^15] https://huggingface.co/datasets/ltg/norcommonsenseqa Mikhailov et al. (2025) consisting of multiple-choice commonsense question answer-pairs which adapts the corresponding English CommonsenseQA dataset Talmor et al. (2019) to Norwegian.
@@
    Summarization, which measures the ability of a model to condense longer pieces of text into shorter, coherent summaries that capture the main points. This skill is crucial for applications where users need a quick understanding of large volumes of information, such as news articles or research papers. It was evaluated using the NorSumm dataset [^17] https://huggingface.co/datasets/SamiaT/NorSumm Touileb et al. (2025).
@@
    Translation, which assesses how accurately a language model can convert a text from one language to another while preserving the meaning, tone, and context. It was evaluated using the existing Tatoeba dataset Tiedemann (2020). The following six language pairs are considered: Bokmål $\leftrightarrow$ Nynorsk, Bokmål $\leftrightarrow$ English, and English $\leftrightarrow$ Nynorsk.
@@
We follow the standard in-context learning evaluation design for pretrained decoder-only language models (e.g., Brown et al., 2020; Touvron et al., 2023), which includes zero-shot and few-shot evaluation. In this paper, for the sake of simplicity, we selected the most common metrics per task and aggregated scores using a simple cumulative sum per higher-level skill. In order to aggregate results into an overall score, with the caveats of aggregating metrics of different nature, scores were extracted for the best available {0, 1, 4, 16}-shot configuration for each task and the best score for each of the prompts. Metrics were normalized to exhibit the same higher-is-better behavior in a range of 0 to 100.
Mistral 7B v0.1base (warm)extended (warm)base + translated booksbase + original books + newspapersbase + original booksbase + nonfiction books + newspapersbase + nonfiction booksbase + fiction booksbase + books + newspapersbase + newspapersbase + booksbaseextended441.83413.98427.30440.35435.64408.20427.40440.97427.87436.81409.54482.98480.28448.01Sentiment AnalysisFairness & TruthfulnessReading ComprehensionWorld KnowledgeCommonsense ReasoningNorwegian LanguageSummarizationTranslationVariation & Readability